OpenStudy (anonymous):

Solve the system by the method of your choice. Identify systems with no solution and systems with infinitely many solutions, using set notation to express their solution sets. 4x - 3y = 6 -12x + 9y = -24 A. {( 3, 4)} B. C. {(x, y) | 4x - 3y = 6 } D. ∅

OpenStudy (anonymous):

Answer is D -12x+9y =-24 can also be written as 3(-4x+3y) = -8*3 This reduces the equation to -4x+3y=-8 or 4x-3y=8 Note the first equation and this equation are nearly the same except the constant term. This implies, The given system is inconsistent, they have no solution. Answer is D

OpenStudy (anonymous):

thank you now i can use this for feature equations.